OpenStudy (anonymous):

Evaluate the integral or show that it is divergent \[\int\limits_{1}^{e} dx / x \sqrt{lnx}\]

myininaya (myininaya):

ok so me know its improper \[\int\limits_{}^{}\frac{dx}{x \sqrt{lnx}},u=lnx=>du=\frac{1}{x} dx\] \[\int\limits_{}^{}\frac{1}{\sqrt{u}} du=\int\limits_{}^{}u^\frac{-1}{2} du\]

myininaya (myininaya):

\[=\frac{u^{\frac{-1}{2}+1}}{\frac{-1}{2}+1} +C\]

myininaya (myininaya):

so so lets get back to the improper integral

OpenStudy (zarkon):

2u^{1/2}

myininaya (myininaya):

you are right

myininaya (myininaya):

forgot to flip

myininaya (myininaya):

\[2u^\frac{1}{2}+C=2(lnx)^\frac{1}{2}+C\] \[\lim_{n \rightarrow 1^+} (2(lnx)^\frac{1}{2})|^e_n\]

myininaya (myininaya):

\[2\lim_{n \rightarrow 1^+}[(\ln(e))^\frac{1}{2}-(\ln(n))^\frac{1}{2}]\]

myininaya (myininaya):

\[2-2\lim_{n \rightarrow 1^+ }(\sqrt{lnx})\]

myininaya (myininaya):

we can use direct substitution :)

myininaya (myininaya):

=2

OpenStudy (zarkon):

yep

myininaya (myininaya):

its getting late

OpenStudy (zarkon):

yep...bed time for me

myininaya (myininaya):

me too goodnight like doing this when i have a partner lol

myininaya (myininaya):

or more i like doing it more when i have a partner

OpenStudy (zarkon):

yep

OpenStudy (zarkon):

night

myininaya (myininaya):

peace